OpenStudy (anonymous):

The radius of a circle is 6 1/4.What is the circumference in terms of pi.

jimthompson5910 (jim_thompson5910):

C = 2*pi*r C = 2*pi*(6 1/4) C = 2*pi*(25/4) .... Note: 6 1/4 = 6+1/4 = 24/4+1/4 = 25/4 C = (2*25/4)*pi C = (50/4)*pi C = (25/2)*pi So the circumference is \[\Large \frac{25}{2}\pi\]

OpenStudy (kymber):